# LLMs Do Not Think Step-by-step In Implicit Reasoning

### v1

@@ 1 Introduction
In conclusion, we think LLMs, despite they can often directly give the correct answer of a multi-step problem, especially when with a larger size, they are not really doing step-by-step reasoning (at least in arithmetic problems), unless adopting explicit CoT. Implicit reasoning may just be an illusion created by LLMs’ powerful memory and rich experience, which is fundamentally different from conventional reasoning. Our study provides critical insights into the mechanics of implicit reasoning and emphasizes the ongoing necessity for explicit CoT methodologies in enhancing LLMs ability on complex tasks.
@@ 2.2 Results of Probing Intermediate Steps
The results in Figure 2 and Figure 3 shows the accuracy of probing the intermediate result of each step when the problem is 3-step or 5-step. It is clear that the results of the first step and the last step can always be probed successfully in the back layers, indicating the model does memorize the input value (i.e. the result of the first step) and does conceive the final answer (i.e. the result of the last step). By contrast, the result of the second step can barely be probed with a lower accuracy, and the results of other steps in the middle can hardly be detected. It looks that the curve of the last step just surges in the last layers, even without waiting for the processing of the 3rd or 4th step.
However, since the result of the second step can be detected to some extent, both in 3-step or 5-step problems, this suggests that LLMs may have the ability to perform a 2-hop reasoning (the 3-step problem actually only needs 2 hops because the result of the first step is already given) in implicit reasoning, but not at all when there are more steps involved.
This finding indicates that, in generic cases, there is actually not a specific state where the model calculates the results of the intermediate steps, even when it correctly give a answer of the multi-step problem. It actually skips the intermediate steps and come up with the final result directly. Therefore, we posit that perhaps due to a large model’s strong abstraction and memory abilities, it has learned a large number of answers to mathematical problems during the training stage. Therefore, it can almost directly map problems with multiple steps to their answers through intuition and memory, thus producing a “implicit reasoning” effect. But in fact, its mechanism is not equivalent to the explicit CoT process at all.
2.3 Result of Slightly Perturbing the Prompt
To further show the difference between implicit reasoning and explicit reasoning, we slightly modify the problem by 2 methods: 1. reversing the order of the equations; 2. dividing all values by 10. For humans, such modifications hardly increase any difficulty of the problem. The examples of the modified problem are as follows:
@@ 2.3 Result of Slightly Perturbing the Prompt
We evaluate the model’s performance in these 3 types of prompts while the original values of the problems are the same. We test with prompt styles of both the implicit reasoning way (as shown in section 2.1) and the explicit reasoning way (adding “let’s think step by step” ).

### v2

@@ 1 Introduction
In conclusion, we think, despite LLMs can often directly give the correct answer of a multi-step problem, especially those with very large sizes, they are not really doing step-by-step reasoning (at least in arithmetic problems), unless adopting explicit CoT. Implicit reasoning may just be an illusion created by LLMs’ powerful memory and rich experience, which is fundamentally different from conventional reasoning. Our study provides critical insights into the mechanics of implicit reasoning and emphasizes the ongoing necessity for explicit CoT methodologies in enhancing LLMs ability on complex tasks.
@@ 2.2 Results of Probing Intermediate Steps
The results in Figure 2 and Figure 3 shows the accuracy of probing the intermediate result of each step when the problem is 3-step or 5-step. It is clear that, in both situations, the results of the first step and the last step can always be probed successfully in the back layers, indicating the model does memorize the input value (i.e. the result of the first step) and does conceive the final answer (i.e. the result of the last step). And the result of the second step can also be detected to some extent, which suggests that LLMs may indeed have the ability to perform a 2-hop reasoning (the 3-step problem actually only needs 2 hops because the result of the first step is already given) in implicit reasoning. This phenomenon is consistent with the conclusion of previous research Yang et al. .
However, by contrast, in the 5-step problem, the results of other steps (3rd and 4th) can hardly be detected. It looks that the curve of the last step just surges in the last layers, even without waiting for the processing of the 3rd or 4th step. Therefore, the model may not calculate the results of 3rd and 4th steps at all.
This finding indicates that, in generic cases where the problem is more than 2-hop, there is actually not a specific state where the model calculates the results of the intermediate steps, even though it can give the correct answer of the multi-step problem. It actually skips the intermediate steps and come up with the final result directly. Therefore, we posit that perhaps due to a large model’s strong abstraction and memory abilities, it has learned a large number of answers to mathematical problems during the training stage. Therefore, it can almost directly map problems with multiple steps to their answers through intuition and experience, thus producing an “implicit reasoning” effect. But in fact, its mechanism is not equivalent to the explicit CoT process at all.
2.3 Result of Modifying the Problem Presentation
To further show the difference in mechanism between implicit reasoning and explicit reasoning, we slightly modify the problem by 2 methods: 1. reversing the order of the equations; 2. dividing all values by 10. Thus we obtain 3 types of problem presentations. The examples of the modified problem are as follows:
@@ 2.3 Result of Modifying the Problem Presentation
For humans or LLMs performing CoT, such modifications hardly increase any difficulty of the problem, because the reasoning steps are not changed at all. However, for a model which only relies on intuition and experience, they will result in a significant difference between the form of the given problem and the form in experience. Through this comparison, we can be more convinced of the difference in mechanism between implicit and explicit reasoning.
We evaluate the model’s performance in under the 3 types of problem presentations while keeping the original values of the problems the same. We test with prompt styles of both the implicit reasoning way (as shown in section 2.1) and the explicit reasoning way (adding “let’s think step by step” ).
@@ 3 Conclusion
4 Limitations
We only test one type of reasoning problems. However, for more complex problems, there may be multiple situations of the intermediate steps and the final answer, making the experiment hard to design.
The fact that the intermediate results cannot be detected by linear probing cannot completely confirm they do not exist. Maybe they are encoded in another form, which is different from that of the results of the first and last step, and is more difficult to be detected. However, we have not yet found a better detection method.